# Scalable Graph Condensation with Evolving Capabilities

### v1

@@ 4.1. A Deep Dive into Condensation Objectives
Intuitively, the condensation process should preserve sufficient information from training data to maintain GNN training performance while ensuring model generalization to test data. Thus, we divide our discussion into two stages: training and test. To simplify our analysis, we adopt the Simplified Graph Convolution (SGC) model as the GNN (Wu et al., 2019), due to its simpler design, similar filtering behavior to GCN (Kipf and Welling, 2016), and its frequent use as a backbone and evaluation model in numerous graph condensation works (Jin et al., 2022b, a; Gong et al., 2024; Xiao et al., 2024).
@@ 4.2. Efficient Optimization for the New Objective
Revolving around the new condensation objective we propose, we now discuss its efficient optimization and the establishment of a correspondence between original and condensed nodes to enhance traceability. As our objective includes representation and parameter distance, we separate their discussions as follows.
@@
By solving the above optimization problem, we can derive the assignment matrix ${\bf P}$ and the propagated features ${\bf F}^{\prime}$ for graph $G^{\prime}$. In alignment with the popular structure-free GC paradigm (Zhang et al., 2024; Zheng et al., 2024a; Jin et al., 2022b), we set $\mathbf{F}^{\prime}$ as the condensed node features and use an identity matrix $\mathbf{I}$ for the condensed adjacency matrix to form a condensed graph that effectively minimizes the representation distance. Notably, the assignment matrix $\mathbf{P}$ and the condensed features $\mathbf{F}^{\prime}$ can be efficiently derived using any Expectation-Maximization (EM)-based clustering algorithm, such as $k$-means. Since $k$-means inherently minimizes the Sum of Squared Errors (SSE), this optimization aligns with the objective $\min\|\mathbf{P}\mathbf{C}-\mathbf{F}\|$, where $\mathbf{C}$ represents the cluster centers of feature points $\mathbf{F}$. The procedure iteratively updates the cluster centroids $\mathbf{F}^{\prime}$ and the assignment matrix $\mathbf{P}$ until convergence is reached. This approach presents two major advantages: (a) Traceability: It provides clear insight into how each original node contributes to the condensed graph, facilitating a better understanding of the condensation process. With traceability, users can manipulate the condensed graph alongside the original one—for example, easily filtering low-quality data when the corresponding original nodes exhibit poor characteristics. (b) Efficiency and Versatility: Unlike traditional GC methods that need specific GNN models to be trained, this new method does not require any GNN training or gradient calculation and is not limited to a particular GNN architecture, which greatly improves efficiency.
@@ 4.3.1. Feature Propagation
The goal of feature propagation is to propagate information from a node’s neighbors to provide each node with a richer representation. To enable a GC procedure that eliminates the need for training, we develop a non-parametric feature propagation module (Wu et al., 2019; Ma et al., 2008; Maurya et al., 2021; Hamilton et al., 2017) designed to produce node embeddings $\mathbf{F}_{t}$ in the original graph $G_{t}$. In our work, by following the propagation method in SGC (Wu et al., 2019), the propagated node representations after $k$ steps are computed as: $\mathbf{F}_{t,k}=\hat{\mathbf{A}}_{t}^{k}\mathbf{X}_{t}$, where $\hat{\mathbf{A}}_{t}$ is the normalized adjacency matrix of $G_{t}$, defined as: $\hat{\mathbf{A}}_{t}=\tilde{\mathbf{D}}_{t}^{-\frac{1}{2}}\tilde{\mathbf{A}}_{t}\tilde{\mathbf{D}}_{t}^{-\frac{1}{2}}$, with $\tilde{\mathbf{A}}_{t}=\mathbf{A}_{t}+\mathbf{I}$ being the adjacency matrix with self-loops, and $\tilde{\mathbf{D}}_{t}$ being the degree matrix of $\tilde{\mathbf{A}}_{t}$, where $(\tilde{\mathbf{D}}_{t})_{i,i}=\sum_{j}\tilde{\mathbf{A}}_{t,i,j}$. To combine information from different propagation steps, we adopt a linearization approach. The final representation for each node is given by: $\mathbf{F}_{t}=\sum_{k=0}^{K}\alpha_{k}\mathbf{F}_{t,k}$. This formulation ensures that each node’s representation captures multi-hop neighborhood information, which fuse the structure and feature information. Note that the weights $\alpha$ can be chosen flexibly. In particular, if $\alpha<0$, it effectively introduces a “negative offset” that can capture heterophilic relationships in the representation (Zhu et al., 2021).
@@ 4.3.2. Representation Clustering
Once the node representations $\mathbf{F}_{t}\in\mathbb{R}^{n_{t}\times d}$ are obtained, we discard the original graph structure and partition the node representations into groups via the following clustering technique. Each node in the condensed graph $G^{\prime}_{t}$ belongs to one of $c$ classes. We denote their labels as $\mathbf{y}^{\prime}_{t}\in\{1,2,\dots,c\}^{n^{\prime}_{t}}$ or equivalently via one-hot encoding $\mathbf{Y}^{\prime}_{t}\in\mathbb{R}^{n^{\prime}_{t}\times c}$. Following the approach of GCond (Jin et al., 2022b), we predefine the labels of the condensed nodes so that their class distribution in $\mathbf{y}^{\prime}_{t}$ matches the original distribution in $\mathbf{y}_{t}$.
A Unified Clustering Formulation. By leveraging the node labels and their feature representations $\{\mathbf{F}_{t,k}\,\mid\,k=1,\dots,c\}$, we can apply a variety of clustering methods to partition the nodes. Here, we adopt a unified formulation that accommodates both hard $k$-means and soft $k$-means. We first split the labeled nodes by their class label $k\in\{1,2,\dots,c\}$. Within class $k$, we assign the nodes to $M_{k}$ clusters, where $M_{k}$ is chosen according to a desired reduction rate $r$ [^1] Reduction rate $r$ is defined as (#nodes in synthetic set)/(#nodes in training set).. Let $M=\sum_{k=1}^{c}M_{k}=n_{t,k}\times\ r,$, where $n_{t,k}$ is number of training nodes in class $k$. Next, we formulate the (hard/soft) assignment matrix $\mathbf{P}_{t}\in\mathbb{R}^{\,n_{t}\times M}$. Each row corresponds to one labeled node, and each column corresponds to $M$ clusters across all classes. In more fine-grained notation, if cluster $j$ belongs to class $k$, then the entry $\mathbf{P}_{t,i,j}$ indicates how node $i$ of class $k$ is associated with that cluster. Formally,
@@
In the hard-assignment scenario, each node belongs to exactly one cluster, so $\mathbf{P}_{t,i,j}=1$ if node $i$ is placed in cluster $j$ and 00 otherwise. In the soft-assignment scenario, each node has fractional memberships across the clusters in its class, with the memberships summing to 1.
@@
where $\mathbf{D}_{P_{t}}\in\mathbb{R}^{n_{t}\times n_{t}}$ is a diagonal matrix satisfying $(\mathbf{D}_{P_{t}})_{k,k}=\sum_{i=1}^{n_{t}}\mathbf{P}_{t,i,k}$. Finally, as the propagated features implicitly encode structural information, and current SOTA methods (Zhang et al., 2024; Zheng et al., 2024a) operate on structure-free graphs, we do not explicitly generate new edges. Beyond achieving strong predictive performance, this structure-free design is significantly more efficient than structure-based approaches, which typically require an additional edge-generation process with $\mathcal{O}((n_{t}r)^{2})$ complexity (Gong et al., 2024).
@@ 4.4. Evolving Condensation via Incremental Initialization
First, let the set of already selected centroids at step $t-1$ be $\mathbf{C}_{t-1}=\{\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{k}\}$, where $\mathbf{c}_{i}\in\mathbb{R}^{d}$, and let the new coming node set in new time step is $\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{m}\}$. For the current step $t$, we compute the distance $D(\mathbf{x}_{i},\mathbf{C}_{t-1})$ for each node embedding $\mathbf{x}_{i}\in\mathbb{R}^{d}$ in class $c$ to consider how well the current centroids represent the data points in the feature space and identify regions of the space that require additional centroids. The distance is defined as the minimum Euclidean distance between $\mathbf{x}_{i}$ and any centroid in $\mathbf{C}_{t-1}$:
@@
Once the centroids $\mathbf{C}_{t}$ are obtained through this process, they will be served as the condensed node representations $\mathbf{X}^{\prime}_{t}$ in the synthetic graph $G^{\prime}_{t}$ at time step $t$. By leveraging this probabilistic initialization, the clustering method ensures that new centroids are well-distributed and far from existing ones, facilitating a more effective condensation of graph representations at each time step.
4.5. Complexity analysis
@@ 5. Experiments
To validate the effectiveness of our proposed GECC, we compare it against classic and SOTA baselines in both non-evolving and evolving scenarios. We first detail the experimental setup, including the construction of benchmark datasets and the experimental settings of each method. Next, we present the node classification performance as a measure of condensation results, alongside a comparison of efficiency. Finally, we empirically demonstrate the effectiveness of the feature aggregation module and the importance of the specific incremental initialization design in our method.
@@ 5.1. Experimental Setup
Datasets and Baselines. Following most of the GC papers, we select seven datasets: five transductive datasets, i.e., Citeseer, Cora (Kipf and Welling, 2016), Pubmed (Namata et al., 2012), Ogbn-arxiv, and Ogbn-products (Hu et al., 2020) and two inductive datasets, Flickr and Reddit (Zeng et al., 2019). All training graphs are randomly divided into five subsets, each preserving the original class distribution. For transductive graphs, nodes in training set are split, whereas inductive graphs are partitioned into subgraphs. Subsequently, the training sets are incrementally enlarged—for example, the first subset forms $G_{1}$, and the first plus the second subset forms $G_{2}$ as formulated in Section 3. For additional dataset details, please refer to Appendix C.1. We compare GECC with most effective and efficient GC methods, encompassing a diverse range of optimization strategies: (1) gradient matching-based: GCond and GCondX (Jin et al., 2022b); (2) distribution matching-based: GCDM (Liu et al., 2022) and SimGC (Xiao et al., 2024); (3) trajectory matching-based: GEOM (Zhang et al., 2024). We also include node selection baselines (Random (Jin et al., 2022b), KCenter (Sener and Savarese, 2017), Herding (Welling, 2009)) to better compare performance–efficiency trade-offs. The condensed graphs are evaluated using a standard GCN trained for 300 epochs with a learning rate of 0.01, as suggested by GC4NC (Gong et al., 2024). The GCN is trained on the synthetic dataset and validated and tested on the original validation and test sets. We also list the results from training a standard GCN in whole dataset for both two settings to show the potential upper bound for graph condensation.
Table 2. Comparison of different condensation methods in two settings and seven datasets. The evolving setting row show the average test accuracy of five time steps. The best results are in bold and the second-best are underlined. ”OOM” indicates out-of-memory errors. Average condensation time (seconds) for both settings are listed alongside test accuracy for each method, with the best values highlighted in bold.
@@
Ogbn-arxiv Non-Evolving 60.19 57.70 58.66 59.93 2895 64.23 6076 60.71 686 65.26 2687 69.59 1685 66.71 10 70.95
@@ 5.2.1. Performance Comparison
To compare the effectiveness of GECC with the baselines, we utilize the generated condensed graph data to train a standard GCN, reporting both test accuracies and standard deviations in Table 2. For each dataset, we present two test accuracies: one for Non-Evolving setting, reflects the test accuracy achieved by training the GCN on the graph at the final time step, $t=T_{5}$, corresponding to the largest possible graph size. Additionally, Figure 1 demonstrates the test accuracy across different time steps which compares the GECC evolving capability with the existing baselines. one from the Evolving setting, which represents the average test accuracy across five time-steps as the graph evolves. This demonstrates the dynamic performance of the condensation method as the graph grows. Our analysis reveals several key insights:
Non-evolving setting – GECC outperforms the baselines on almost all datasets, including the whole dataset performance for some datasets, which highlights the effectiveness of our training-free approach in the non-evolving setting. The only exceptions are observed in the Flickr and Ogbn-arxiv datasets, where GECC ranks second by a minimal margin. In addition, unlike other baselines such as GCond, GECC does not fail on extremely large datasets like Ogbn-products due to memory constraints.
Evolving setting – By analyzing Table  2, we observe that GECC consistently outperforms existing baselines across almost all datasets, often by a large margin. The only exception is Ogbn-arxiv, where GECC ranks second. However, Figure 1 highlights that GECC achieves nearly its maximum possible test accuracy even at very early time steps. For instance, on Ogbn-arxiv, it surpasses 65% accuracy by the second time step, a feat unmatched by existing baselines. This demonstrates GECC’s efficiency in leveraging limited data for superior generalization, whereas other methods struggle to reach comparable performance early in the training process. Moreover, we observe that GECC surpasses the performance achieved on the whole dataset for relatively smaller graphs such as Cora, Citeseer, and Pubmed. This result highlights the effectiveness of our training-free condensation approach in achieving a ”lossless” objective during graph evolution. Additionally, Figure 1 illustrates the robustness of GECC in steadily improving performance as training size increases. This trend fails to hold for baselines, particularly for GEOM on Reddit. We conjecture that the trajectory matching method heavily relies on the performance of the pre-trained GNN, making it susceptible to the twofold influence of data size changes: first, during the pre-training stage, and second, during the condensation stage.

### v2

@@ 4.1. A Deep Dive into Condensation Objectives
Intuitively, the condensation process should preserve sufficient information from training data to maintain GNN training performance while ensuring model generalization to test data. Thus, we divide our discussion into two stages: training and test. To simplify our analysis, we adopt the Simplified Graph Convolution (SGC) model as the GNN (Wu et al., 2019), due to its simpler design, similar filtering behavior to GCN (Kipf and Welling, 2016), and its frequent use as a backbone and evaluation model in numerous graph condensation works (Jin et al., 2021, 2022; Gong et al., 2024; Xiao et al., 2024).
@@ 4.2. Efficient Optimization for the New Objective
Focusing on the new condensation objective we propose, we now discuss its efficient optimization and the establishment of a correspondence between original and condensed nodes to enhance traceability. As our objective includes representation and parameter distance, we separate their discussions as follows.
@@
By solving the above optimization problem, we can derive the assignment matrix ${\bf P}$ and the propagated features ${\bf F}^{\prime}$ for graph $G^{\prime}$. In alignment with the popular structure-free GC paradigm (Zhang et al., 2024; Zheng et al., 2024; Jin et al., 2021), we set $\mathbf{F}^{\prime}$ as the condensed node features and use an identity matrix $\mathbf{I}$ for the condensed adjacency matrix to form a condensed graph that effectively minimizes the representation distance. Notably, the assignment matrix $\mathbf{P}$ and the condensed features $\mathbf{F}^{\prime}$ can be efficiently derived using any Expectation-Maximization (EM)-based clustering algorithm, such as $k$-means. Since $k$-means inherently minimizes the Sum of Squared Errors (SSE), this optimization aligns with the objective $\min\|\mathbf{P}\mathbf{C}-\mathbf{F}\|$, where $\mathbf{C}$ represents the cluster centers of feature points $\mathbf{F}$. The procedure iteratively updates the cluster centroids $\mathbf{F}^{\prime}$ and the assignment matrix $\mathbf{P}$ until convergence is reached. This approach presents two major advantages: (a) Traceability: It provides clear insight into how each original node contributes to the condensed graph, facilitating a better understanding of the condensation process. With traceability, users can manipulate the condensed graph alongside the original one—for example, easily filtering low-quality data when the corresponding original nodes exhibit poor characteristics. (b) Efficiency and Versatility: Unlike traditional GC methods that need specific GNN models to be trained, this new method does not require any GNN training or gradient calculation and is not limited to a particular GNN architecture, which greatly improves efficiency.
@@ 4.3.1. Feature Propagation
The goal of feature propagation is to propagate information from a node’s neighbors to provide each node with a richer representation. To enable a GC procedure that eliminates the need for training on the whole graphs, we develop a non-parametric feature propagation module (Wu et al., 2019; Ma et al., 2008; Maurya et al., 2021; Hamilton et al., 2017) designed to produce node embeddings $\mathbf{F}_{t}$ in the original graph $G_{t}$. In our work, by following the propagation method in SGC (Wu et al., 2019), the propagated node representations after $k$ steps are computed as: $\mathbf{F}_{t,k}=\hat{\mathbf{A}}_{t}^{k}\mathbf{X}_{t}$, where $\hat{\mathbf{A}}_{t}$ is the normalized adjacency matrix of $G_{t}$, defined as: $\hat{\mathbf{A}}_{t}=\tilde{\mathbf{D}}_{t}^{-\frac{1}{2}}\tilde{\mathbf{A}}_{t}\tilde{\mathbf{D}}_{t}^{-\frac{1}{2}}$, with $\tilde{\mathbf{A}}_{t}=\mathbf{A}_{t}+\mathbf{I}$ being the adjacency matrix with self-loops, and $\tilde{\mathbf{D}}_{t}$ being the degree matrix of $\tilde{\mathbf{A}}_{t}$, where $(\tilde{\mathbf{D}}_{t})_{i,i}=\sum_{j}\tilde{\mathbf{A}}_{t,i,j}$. To combine information from different propagation steps, we adopt a linearization approach. The final representation for each node is given by: $\mathbf{F}_{t}=\sum_{k=0}^{K}\alpha_{k}\mathbf{F}_{t,k}$. This formulation ensures that each node’s representation captures multi-hop neighborhood information, which fuse the structure and feature information. Note that we allow the weights $\alpha$ to vary in a large range. In particular, if $\alpha<0$, it effectively introduces a “negative offset” that can capture heterophilic relationships in the representation (Zhu et al., 2021).
@@ 4.3.2. Representation Clustering
Once the node representations $\mathbf{F}_{t}\in\mathbb{R}^{n_{t}\times d}$ are obtained, we discard the original graph structure and partition the node representations into groups via the following clustering technique. Each node in the condensed graph $G^{\prime}_{t}$ belongs to one of $c$ classes. We denote their labels as $\mathbf{y}^{\prime}_{t}\in\{1,2,\dots,c\}^{n^{\prime}_{t}}$ or equivalently via one-hot encoding $\mathbf{Y}^{\prime}_{t}\in\mathbb{R}^{n^{\prime}_{t}\times c}$. Following the approach of GCond (Jin et al., 2021), we predefine the labels of the condensed nodes so that their class distribution in $\mathbf{y}^{\prime}_{t}$ matches the original distribution in $\mathbf{y}_{t}$.
A Unified Clustering Formulation. By leveraging the node labels and their feature representations $\{\mathbf{F}_{t,k}\,\mid\,k=1,\dots,c\}$, we can apply a variety of clustering methods to partition the nodes. Here, we adopt a unified formulation that accommodates both hard $k$-means and soft $k$-means. We first split the labeled nodes by their class label $k\in\{1,2,\dots,c\}$. Within class $k$, we assign the nodes to $M_{k}$ clusters, where $M_{k}$ is chosen according to a desired reduction rate $r$ [^1] Reduction rate $r$ is defined as (#nodes in synthetic set)/(#nodes in training set).. Let $M=\sum_{k=1}^{c}M_{k}=n_{t,k}\times\ r$, where $n_{t,k}$ is number of training nodes in class $k$. Next, we formulate the (hard/soft) assignment matrix $\mathbf{P}_{t}\in\mathbb{R}^{\,n_{t}\times M}$. Each row corresponds to one labeled node, and each column corresponds to $M$ clusters across all classes. In more fine-grained notation, if cluster $j$ belongs to class $k$, then the entry $\mathbf{P}_{t,i,j}$ indicates how node $i$ of class $k$ is associated with that cluster. Formally,
@@
In the hard-assignment scenario, each node belongs to exactly one cluster, so $\mathbf{P}_{t,i,j}=1$ if node $i$ is placed in cluster $j$ and 0 otherwise. In the soft-assignment scenario, each node has fractional memberships across the clusters in its class, with the memberships summing to 1.
@@
where $\mathbf{D}_{P_{t}}\in\mathbb{R}^{n_{t}\times n_{t}}$ is a diagonal matrix satisfying $(\mathbf{D}_{P_{t}})_{k,k}=\sum_{i=1}^{n_{t}}\mathbf{P}_{t,i,k}$. Finally, as the propagated features implicitly encode structural information, and current SOTA methods (Zhang et al., 2024; Zheng et al., 2024) operate on structure-free graphs, we do not explicitly generate new edges. Beyond achieving strong predictive performance, this structure-free design is significantly more efficient than structure-based approaches, which typically require an additional edge-generation process with $\mathcal{O}((n_{t}r)^{2})$ complexity (Gong et al., 2024).
@@ 4.4. Evolving Condensation via Incremental Initialization
First, let the set of already selected centroids at step $t-1$ be $\mathbf{C}_{t-1}=\{\mathbf{c}_{1},\mathbf{c}_{2},\dots,\mathbf{c}_{k}\}$, where $\mathbf{c}_{i}\in\mathbb{R}^{d}$, and let the new coming node set in new time step is $\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{m}\}$. For the current step $t$, we compute $D(\mathbf{x}_{i},\mathbf{C}_{t-1})$ for each node embedding $\mathbf{x}_{i}\in\mathbb{R}^{d}$ in class $c$ to consider how well the current centroids represent the data points in the feature space and identify regions of the space that require additional centroids. The distance is defined as the minimum Euclidean distance between $\mathbf{x}_{i}$ and any centroid in $\mathbf{C}_{t-1}$:
@@
Once the centroids $\mathbf{C}_{t}$ are obtained through this process, they will be served as the condensed node representations $\mathbf{X}^{\prime}_{t}$ in the synthetic graph $G^{\prime}_{t}$ at time step $t$. By leveraging this probabilistic initialization, the clustering method ensures that new centroids are well-distributed and far from existing ones, facilitating a more effective condensation of graph representations at each time step. The complexity analysis of the whole method can be found in Appendix D.
@@ 5. Experiments
To validate the effectiveness of our proposed GECC, we compare it against classic and SOTA baselines in both non-evolving and evolving scenarios. We first detail the experimental setup, including the construction of benchmark datasets and the experimental settings of each method. Next, we present the node classification performance as a measure of condensation results, alongside a comparison of efficiency. Finally, we empirically demonstrate the effectiveness of the feature aggregation module and the importance of the specific incremental initialization design in our method. Our code is provided in https://anonymous.4open.science/r/GECC-F4CA.
@@ 5.1. Experimental Setup
Datasets and Baselines. Following most of the GC papers, we select seven datasets: five transductive datasets, i.e., Citeseer, Cora (Kipf and Welling, 2016), Pubmed (Namata et al., 2012), Ogbn-arxiv, and Ogbn-products (Hu et al., 2020) and two inductive datasets, Flickr and Reddit (Zeng et al., 2019). All training graphs are randomly divided into five subsets and each preserving the original class distribution except for Ogbn-arxiv-real, which follows the real timestamps of the publication years. For transductive graphs, nodes in training set are split, whereas inductive graphs are partitioned into subgraphs. Subsequently, the training sets are incrementally enlarged—for example, the first subset forms $G_{1}$, and the first plus the second subset forms $G_{2}$. For additional dataset details, please refer to Appendix E.1. We compare GECC with most effective and efficient GC methods, encompassing a diverse range of optimization strategies: (1) gradient matching-based: GCond and GCondX (Jin et al., 2021); (2) distribution matching-based: GCDM (Liu et al., 2022) and SimGC (Xiao et al., 2024); (3) trajectory matching-based: GEOM (Zhang et al., 2024). We also include node selection baselines (Random (Jin et al., 2021), KCenter (Sener and Savarese, 2017), Herding (Welling, 2009)) to better compare performance–efficiency trade-offs. The condensed graphs are evaluated using a standard GCN trained for 300 epochs with a learning rate of 0.01, as suggested by GC4NC (Gong et al., 2024). The GCN is trained on the synthetic dataset and validated and tested on the original validation and test sets. We also list the results from training a standard GCN in whole dataset for both two settings to show the potential upper bound for graph condensation.
Table 2. Comparison of different condensation methods in two settings and seven datasets. The evolving setting row show the average test accuracy of five time steps. The best results are in bold and the second-best are underlined. ”OOM” indicates out-of-memory errors. Average condensation time (seconds) for both settings are listed alongside test accuracy for each method.
Dataset Setting Random Herding KCenter GCondX GCDM GCond SimGC GEOM GECC Whole
@@
Ogbn-arxiv Non-Evolving 60.19 57.70 58.66 59.93 2895 64.23 3020 60.71 686 65.26 2687 69.59 1685 66.71 10 70.95
@@ 5.2.1. Performance Comparison
To compare the effectiveness of GECC with the baselines, we utilize the generated condensed graph data to train a standard GCN, reporting both test accuracies and standard deviations in Table 2. For each dataset, we report two test accuracies: one for the Non-Evolving setting, reflecting the accuracy achieved by training the GCN on the graph at the final time step ($t=5$), corresponding to the largest graph size; and one for the Evolving setting, representing the average accuracy across all five time steps as the graph evolves. It is also noteworthy that, to evaluate our proposed method on real-world dynamic datasets such as Ogbn-Arxiv-real, we utilize the original metadata from Ogbn-arxiv (Hu et al., 2020) and use node timestamps to divide the data into five sequential time steps.
Additionally, Figure 1 illustrates the test accuracy across different time steps, which compares the GECC evolving capability with the existing baselines. Our analysis reveals several key insights:
Non-evolving setting – GECC outperforms the baselines on almost all datasets, including the whole dataset performance for some datasets, which highlights the effectiveness of our training-free approach in the non-evolving setting. The only exceptions are observed in the Flickr and Ogbn-arxiv datasets, where GECC ranks second by a minimal margin. In addition, unlike other baselines, GECC does not fail on extremely large datasets like Ogbn-products due to memory constraints.
Evolving setting – By analyzing Table  2, we observe that GECC consistently outperforms existing baselines across almost all datasets, often by a large margin. The only exception is Ogbn-arxiv, where GECC ranks second. ALso, Figure 1 highlights that GECC achieves the best test accuracy at early time steps. For instance, on Ogbn-arxiv, it surpasses 65% accuracy by the second time step, a large gap compared to existing baselines. This demonstrates GECC’s efficiency in leveraging limited data for superior generalization, whereas other methods struggle to reach comparable performance early because they need sufficient data for training. For the Ogbn-arxiv-real dataset, a real-world graph with inherent dynamics based on original timestamps, GECC still achieves the best evolving performance, indicating that our method remains effective in realistic dynamic scenarios. Moreover, we observe that GECC surpasses the performance achieved on the whole dataset for relatively smaller graphs such as Cora, Citeseer, and Pubmed. This result highlights the effectiveness of our training-free condensation approach in achieving a ”lossless” objective during graph evolution. Additionally, Figure 1 illustrates the robustness of GECC in steadily improving performance as training size increases. This trend fails to hold for baselines, particularly for GEOM on Reddit. We conjecture that the trajectory matching method heavily relies on the performance of the pre-trained GNN, making it susceptible to the twofold influence of data size changes: first, during the pre-training stage, and second, during the condensation stage.